# Union of Finitely Generated Congruences on Ground Term Algebra

### v1

@@ 8 Constructing the Auxiliary dpwpa for the GTESs $E$ and $F$
if $\sigma(b_{1},\ldots,b_{m})\dot{\rightarrow}b\not\in R\langle F,E\rangle$
@@
$R\langle F,E\rangle:=R\langle F,E\rangle\cup\{\,\sigma(b_{1},\ldots,b_{m})\dot{\rightarrow}b\,\}$;
{ Store $R\langle F,E\rangle$ without duplicates in the red-black tree $RBT8$. }
$\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{F}:=\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{F}+1$;
@@
We build the adjacency-list representation of the dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ by reading in search key order the nodes of the black-red tree $RBT6$ storing the relation $BSTEP\langle E\cup F,\emptyset\rangle$. Recall that for every ordered pair $(a,b)\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component $a$, the second component, $b$ is satellite data.
@@
for each $a\in C\langle E\cup F,\emptyset\rangle$ do
$a.keeps_{E}:=\mathrm{true}$;
for each rewrite rule $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle E\cup F,\emptyset\rangle$ with $m\in{\mathbb{N}}$ do
if $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a.counter_{E}\neq\Pi_{i=1}^{m}|a_{i}/_{\Theta\langle E,F\rangle}|$
then $a.keeps_{E}:=\mathrm{false}$
@@
for each $a\in C\langle E\cup F,\emptyset\rangle$ do
$a.keeps_{F}:=\mathrm{true}$;
for each rewrite rule $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle E\cup F,\emptyset\rangle$ with $m\in{\mathbb{N}}$ do
if $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a.counter_{F}\neq\Pi_{i=1}^{m}|a_{i}/_{\Theta\langle F,E\rangle}|$
then $a.keeps_{F}:=\mathrm{false}$
@@ 1. Algorithm CAD on $\Sigma$, $E$, and $F$ produces
  • $BSTEP\langle E\cup F,\emptyset\rangle$ stored without duplicate values in a red-black tree saved in a variable $RBT6$; for every ordered pair$(a,b)\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component $a$, the second component, $b$ is satellite data;
@@
2. Algorithm CAD on $\Sigma$, $E$, and $F$ decides for every constant $a\in C\langle E\cup F,\emptyset\rangle$, whether $a\in C\langle E,F\rangle$ and whether $a\in C\langle F,E\rangle$, and sets the value of the attribute $a.equal_{E}$ to true if and only if $a\in C\langle E,F\rangle$, and sets the value of the attribute $a.equal_{F}$ to true if and only if $a\in C\langle F,E\rangle$.
[…]
@@
By Proposition 2.16, we know the cardinality of all $W/_{\Theta\langle E,F\rangle}$-classes, $W/_{\Theta\langle F,E\rangle}$-classes, and $W/_{\Theta\langle E\cup F,\emptyset\rangle}$-classes. For all $a\in R\langle E,F\rangle$ and $c\in R\langle E\cup F,\emptyset\rangle$ such that $a\cap c\neq\emptyset$, we have $a=c$ if and only if $|a|=|c|$. Hence we have Statement 2.
We now show Statement 3. For every rewrite rule $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c\in R\langle E\cup F,\emptyset\rangle$, we decide whether for all $a_{1},\ldots,a_{m}\in C\langle E,F\rangle$ with $a_{1}\subseteq c_{1},\ldots,a_{m}\subseteq c_{m}$, there exists a rewrite rule with left-hand side $\sigma(a_{1},\ldots,a_{m})$ in $R\langle E,F\rangle$. When constructing the GTRS $R\langle E,F\rangle$, whenever we add to $R\langle E,F\rangle$ a new rule with left-hand side $\sigma(a_{1},\ldots,a_{m})$, where $a_{1}\subseteq a_{1},\ldots,a_{m}\subseteq c_{m}$, we add $1$ to the attribute $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{E}$. In this way, we count the rules in $R\langle E,F\rangle$ with left-hand side $\sigma(a_{1},\ldots,a_{m})$, where $a_{1}\subseteq c_{1},\ldots,a_{m}\subseteq c_{m}$, and store the result in the attribute $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{E}$. Then $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{E}=\Pi_{i=1}^{m}|c_{i}/_{\Theta\langle E,F\rangle}|$ if and only if for all $a_{1},\ldots,a_{m}\in C\langle E,F\rangle$ with $a_{1}\subseteq c_{1},\ldots,a_{m}\subseteq c_{m}$, there exists a rewrite rule with left-hand side $\sigma(a_{1},\ldots,a_{m})$ in $R\langle E,F\rangle$. When Algorithm CAD terminates, for every vertex $c\in C\langle E\cup F,\emptyset\rangle$,
$c.keeps_{E}=\mathrm{true}$ if and only if
for every rewrite rule of the form $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c\in R\langle{E\cup F},\emptyset\rangle$,
[…]
there exists a rewrite rule in $R\langle E,F\rangle$ of the form $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a$ for some $a\in C\langle E,F\rangle$.
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$. Algorithm CAD runs in $O(n\,\mathrm{log}\,n)$ time on $\Sigma$, $E$, and $F$, where $n=size(E)+size(F)$.
@@
The proof consists of showing the following nine statements:
1. The time complexity of constructing the red-black trees $RBT1$ and $RBT2$ is $O(n)$.
2. The time complexity of constructing the relation $\tau\langle E\cup F,\emptyset\rangle$ and the GTES $E\cup F$ stored without duplicate values in a red-black tree saved in the variable $RBT5$, is $O(n^{2})$ time.
3. The time complexity of computing the value attributes $a.equal_{E}$ and $a.equal_{F}$ for all constants $a\in C\langle E\cup F,\emptyset\rangle$, is $O(n)$.
4. The time complexity of constructing the following data is $O(n\,\mathrm{log}\,n)$.
@@
5. The time complexity of constructing
@@
6. The time complexitity of computing the values of the satellite data $counter_{E}$, and $counter_{F}$ for all rules in $R\langle E\cup F,\emptyset\rangle$ is $O(n)$.
7. The time complexity of constructing the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ is $O(n)$.
8. The time complexitity of computing the values of the attributes $equal_{E}$ and $equal_{F}$ of all constants $a\in C\langle E\cup F,\emptyset\rangle$ is $O(n)$.
9. The time complexitity of computing the values of the attributes $keeps_{E}$ and $keeps_{F}$ of all constants $a\in C\langle E\cup F,\emptyset\rangle$ is $O(n\,\mathrm{log}\,n)$ time.
First we now show Statement 1. By Statement 6.59, $INC(E\cup F,E)leqn$. Consequently, when we construct the red-black tree $RBT1$, we carry out at most $n$ insertion. By Proposition 2.2, we construct the red-black tree $RBT1$ in $O(n\,\mathrm{log}\,n)$ time. Similarly, we construct the red-black tree $RBT2$ in $O(n\,\mathrm{log}\,n)$ time.
Second, by Statement 2.4, we have Statement 2.
Third, we carry out $2n$ comparisons of nonnegative numbers in $O(n)$ time.
Fourth, By Statement 2 and Statement 3 of 4.47, we have Statement 4.
Fifth, we now show Statement 5. When we construct the red-black tree $RBT3$, we read the whole red-black tree $RBT1$ in search key order in $O(n)$ time. Each modification takes constant time. By Statement 6.59, $|INC(E\cup F,E)|\leq n$. Consequently, we construct the red-black tree $RBT3$ in $O(n)$ time. Similarly, we construct the red-black tree $RBT4$ in $O(n)$ time.
Sixth, we now show Statement 6. When Algorithm CAD computes the values of the satellite data $counter_{E}$ and $counter_{F}$ for all rules of the GTRS $R\langle E\cup F,\emptyset\rangle$, it carries out at most $2n$ additions, since $|C\langle E,F\rangle|\leq|W|\leq n$ and $|C\langle F,E\rangle|\leq|W|\leq n$.
Seventh we now show Statement 7. We build the adjacency-list representation of the dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ by reading in search key order the nodes of the black-red tree $RBT6$ storing the relation $BSTEP\langle E\cup F,\emptyset\rangle$. Recall that for every ordered pair $(a,b)\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component $a$, the second component, $b$ is satellite data. Then the time complexity of building the adjacency-list representation of the dpwpa $AUX[E;F]$ is $O(|C\langle E\cup F,\emptyset\rangle|+|A[E;F]|)$. By Statement 4.27 and Statement 4.45 $|C\langle E\cup F,\emptyset\rangle|\leq|W|\leq n$. By Statement 6.63, $|A[E;F]|=|BSTEP\langle E\cup F,\emptyset\rangle|\leq size(E)+size(F)=n$. Therefore the time complexity of building the adjacency-list representation of the dpwpa $AUX[E;F]$ is $O(n)$.
Eighth, Statement 8.78 implies Statement 8.
Nineth, we now show Statement 9. We now estimate the number of operations when computing the values of the attributes $a.keeps_{E}$ for all $a\in C\langle E\cup F,\emptyset\rangle$. For each rewrite rule $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle E\cup F,\emptyset\rangle$, Algorithm CAD carries out $m$ searches in the red-black tree $RBT3$ and less than or equal to $m$ multiplications. By Statement 4.45, $size(R\langle E\cup F,\emptyset\rangle)\leq 2\cdot n$. In this way, when we compute the values of the attribute $keeps_{E}$ for all rewrite rules $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle E\cup F,\emptyset\rangle$, the number of searches in the red-black tree $RBT3$ is less than equal to $n$. Consequently, by Proposition 2.1, the searches in the red-black tree $RBT3$ take $O(n\,\mathrm{log}\,n)$ time.
Furthermore,
the number of multiplications $\leq\sum_{\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle E\cup F,\emptyset\rangle}m\leq size(R\langle E\cup F,\emptyset\rangle)\leq 2n$
by Statement 4.45. Thus, the number of multiplications is less than or equal to $n$.
For each $a\in C\langle E\cup F,\emptyset\rangle$, the number of comparisons of nonnegative integers during computing the value of the attribute $a.keeps_{E}$ is equal to the number of rules in $R\langle E\cup F,\emptyset\rangle$ with right-hand side $a$. Consequently, the number of comparisons of nonnegative integers during computing the values of the attributes $a.keeps_{E}$ for all $a\in C\langle E\cup F,\emptyset\rangle$ is equal to $|R\langle E\cup F,\emptyset\rangle|$. By Statement 4.45
$|R\langle E\cup F,\emptyset\rangle|\leq n$.
Consequently the time complexity of computing for all $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attribute $a.keeps_{E}$ is $O(n\,\mathrm{log}\,n)$. Similarly, the time complexity of computing for all $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attribute $a.keeps_{F}$ is $O(n\,\mathrm{log}\,n)$.
The Lemma follows from Statements 1, 2, 3, 4, 5, 6, 7, 8, and 9. ∎
Statement 8.78.
1. We can decide in $O(n)$ time whether for all $a\in C\langle E\cup F,\emptyset\rangle$, $a\in C\langle E,F\rangle$.
2. We can decide in $O(n)$ time whether for all $a\in C\langle E\cup F,\emptyset\rangle$, $a\in C\langle F,E\rangle$.
@@
First we show Statement 1. We read each nodes of the red-black tree $RBT3$ in lexicographic order. For every $a\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $(a,k)\in NUM(E\cup F,E)$. If $k=1$, then by Statement 6.60, $a\in C\langle E,F\rangle$. Otherwise, $a\not\in C\langle E,F\rangle$.
We can show Statement 2 similarly to Statement 1. ∎

### v2

@@ 8 Constructing the Auxiliary dpwpa for the GTESs $E$ and $F$
if $\sigma({c}_{1},\ldots,{c}_{m})\dot{\rightarrow}{c}\not\in R\langle F,E\rangle$
@@
$R\langle F,E\rangle:=R\langle F,E\rangle\cup\{\,\sigma({c}_{1},\ldots,{c}_{m})\dot{\rightarrow}{c}\,\}$;
(Store $R\langle F,E\rangle$ without duplicates in the red-black tree $RBT8$.)
$\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}.counter_{F}:=\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}.counter_{F}+1$;
@@
We build the adjacency-list representation of the dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ by reading in search key order the nodes of the black-red tree $RBT6$ storing the relation $BSTEP\langle E\cup F,\emptyset\rangle$. Recall that for every ordered pair $({b},{c})\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component b{b}, the second component, c{c} is satellite data.
@@
for each ${b}\in C\langle E\cup F,\emptyset\rangle$ do
${b}.keeps_{E}:=\mathrm{true}$;
for each rewrite rule $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\in R\langle E\cup F,\emptyset\rangle$ with $m\in{\mathbb{N}}$ do
if $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}.counter_{E}\neq\Pi_{i=1}^{m}|{b}_{i}/_{\Theta\langle E,F\rangle}|$
then ${b}.keeps_{E}:=\mathrm{false}$
@@
for each ${b}\in C\langle E\cup F,\emptyset\rangle$ do
${b}.keeps_{F}:=\mathrm{true}$;
for each rewrite rule $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\in R\langle E\cup F,\emptyset\rangle$ with $m\in{\mathbb{N}}$ do
if $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}.counter_{F}\neq\Pi_{i=1}^{m}|{b}_{i}/_{\Theta\langle F,E\rangle}|$
then ${b}.keeps_{F}:=\mathrm{false}$
@@ 1. Algorithm CAD on $\Sigma$, $E$, and $F$ produces
  • 8.
[…]
  • 9.
@@
2. Algorithm CAD on $\Sigma$, $E$, and $F$ decides for every constant ${b}\in C\langle E\cup F,\emptyset\rangle$, whether ${b}\in C\langle E,F\rangle$ and whether ${b}\in C\langle F,E\rangle$, and sets the value of the attribute ${b}.equal_{E}$ to true if and only if ${b}\in C\langle E,F\rangle$, and sets the value of the attribute ${b}.equal_{F}$ to true if and only if ${b}\in C\langle F,E\rangle$.
3. Algorithm CAD on $\Sigma$, $E$, and $F$ decides for any ${d}\in R\langle{E\cup F},\emptyset\rangle$ and any rewrite rule $\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}\in R\langle{E\cup F},\emptyset\rangle$ with right-hand side d{d}, whether for all ${b}_{1},\ldots,{b}_{m}\in C\langle E,F\rangle$ such that ${b}_{i}\subseteq{d}_{i}$ for every $i\in\{\,1,\ldots,m\,\}$, there exists a rewrite rule in $R\langle E,F\rangle$ with left-hand side $\sigma({b}_{1},\ldots,{b}_{m})$. If so, then it sets the value of the attribute ${d}.keeps_{E}$ to $\mathrm{true}$, otherwise it sets the value of the attribute ${d}.keeps_{E}$ to $\mathrm{false}$.
4. Algorithm CAD on $\Sigma$, $E$, and $F$ decides for any ${d}\in R\langle{E\cup F},\emptyset\rangle$ and any rewrite rule $\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}\in R\langle{E\cup F},\emptyset\rangle$ with right-hand side d{d}, whether for all ${c}_{1},\ldots,{c}_{m}\in C\langle F,E\rangle$ such that ${c}_{i}\subseteq{d}_{i}$ for every $i\in\{\,1,\ldots,m\,\}$, there exists a rewrite rule in $R\langle F,E\rangle$ with left-hand side $\sigma({c}_{1},\ldots,{c}_{m})$. If so, then it sets the value of the attribute ${d}.keeps_{F}$ to $\mathrm{true}$, otherwise it sets the value of the attribute ${d}.keeps_{F}$ to $\mathrm{false}$.
@@
By Proposition 2.17, we know the cardinality of all $W/_{\Theta\langle E,F\rangle}$-classes, $W/_{\Theta\langle F,E\rangle}$-classes, and $W/_{\Theta\langle E\cup F,\emptyset\rangle}$-classes. For all ${b}\in R\langle E,F\rangle$ and ${d}\in R\langle E\cup F,\emptyset\rangle$ such that ${b}\cap{d}\neq\emptyset$, we have b=d{b}={d} if and only if |b|=|d||{b}|=|{d}|. Hence we have Statement 2.
We now show Statement 3. For every rewrite rule $\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}\in R\langle E\cup F,\emptyset\rangle$, we decide whether for all ${b}_{1},\ldots,{b}_{m}\in C\langle E,F\rangle$ with ${b}_{1}\subseteq{d}_{1},\ldots,{b}_{m}\subseteq{d}_{m}$, there exists a rewrite rule with left-hand side $\sigma({b}_{1},\ldots,{b}_{m})$ in $R\langle E,F\rangle$. When constructing the GTRS $R\langle E,F\rangle$, whenever we add to $R\langle E,F\rangle$ a new rule with left-hand side $\sigma({b}_{1},\ldots,{b}_{m})$, where ${b}_{1}\subseteq{b}_{1},\ldots,{b}_{m}\subseteq{d}_{m}$, we add $1$ to the attribute $\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}.counter_{E}$. In this way, we count the rules in $R\langle E,F\rangle$ with left-hand side $\sigma({b}_{1},\ldots,{b}_{m})$, where ${b}_{1}\subseteq{d}_{1},\ldots,{b}_{m}\subseteq{d}_{m}$, and store the result in the attribute $\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}.counter_{E}$. Then $\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}.counter_{E}=\Pi_{i=1}^{m}|{d}_{i}/_{\Theta\langle E,F\rangle}|$ if and only if for all ${b}_{1},\ldots,{b}_{m}\in C\langle E,F\rangle$ with ${b}_{1}\subseteq{d}_{1},\ldots,{b}_{m}\subseteq{d}_{m}$, there exists a rewrite rule with left-hand side $\sigma({b}_{1},\ldots,{b}_{m})$ in $R\langle E,F\rangle$. When Algorithm CAD terminates, for every vertex ${d}\in C\langle E\cup F,\emptyset\rangle$,
${d}.keeps_{E}=\mathrm{true}$ if and only if
for every rewrite rule of the form $\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}\in R\langle{E\cup F},\emptyset\rangle$,
for every ${b}_{1},\ldots,{b}_{m}\in C\langle E,F\rangle$ with ${b}_{i}\subseteq{d}_{i}$ for all $i\in\{\,1,\ldots,m\,\}$,
@@
Statement 8.78.
1. We can decide in $O(n)$ time whether for all ${b}\in C\langle E\cup F,\emptyset\rangle$, ${b}\in C\langle E,F\rangle$.
2. We can decide in $O(n)$ time whether for all ${b}\in C\langle E\cup F,\emptyset\rangle$, ${b}\in C\langle F,E\rangle$.
Proof.
First we show Statement 1. We read each nodes of the red-black tree $RBT3$ in lexicographic order. For every ${b}\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $({b},k)\in NUM(E\cup F,E)$. If $k=1$, then by Statement 6.61, ${b}\in C\langle E,F\rangle$. Otherwise, ${b}\not\in C\langle E,F\rangle$.
We can show Statement 2 similarly to Statement 1. ∎
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$. Algorithm CAD runs in $O(n\,\mathrm{log}\,n)$ time on $\Sigma$, $E$, and $F$, where $n=\mathrm{size}(E)+\mathrm{size}(F)$.
@@
We show Statements 8.808.88.
Statement 8.80.
The time complexity of constructing the red-black trees $RBT1$ and $RBT2$ is $O(n)$.
Proof.
By Statement 6.60, $INC(E\cup F,E)\leq n$. Consequently, when we construct the red-black tree $RBT1$, we carry out at most $n$ insertion. By Proposition 2.2, we construct the red-black tree $RBT1$ in $O(n\,\mathrm{log}\,n)$ time. Similarly, we construct the red-black tree $RBT2$ in $O(n\,\mathrm{log}\,n)$ time. ∎
By Statement 2.5, we have the following result.
Statement 8.81.
The time complexity of constructing the relation $\tau\langle E\cup F,\emptyset\rangle$ and the GTES $E\cup F$ stored without duplicate values in a red-black tree saved in the variable $RBT5$, is $O(n^{2})$ time.
We carry out $2n$ comparisons of nonnegative numbers in $O(n)$ time. Hence we have the following fact.
Statement 8.82.
The time complexity of computing the value attributes ${b}.equal_{E}$ and ${b}.equal_{F}$ for all constants ${b}\in C\langle E\cup F,\emptyset\rangle$, is $O(n)$.
By Statement 2 and Statement 3 of 4.48, we have the following fact.
Statement 8.83.
The time complexity of constructing the following data is $O(n\,\mathrm{log}\,n)$.
  • 1.
@@
Statement 8.84.
The time complexity of constructing
  • 1.
@@
When we construct the red-black tree $RBT3$, we read the whole red-black tree $RBT1$ in search key order in $O(n)$ time. Each modification takes constant time. By Statement 6.60, $|INC(E\cup F,E)|\leq n$. Consequently, we construct the red-black tree $RBT3$ in $O(n)$ time. Similarly, we construct the red-black tree $RBT4$ in $O(n)$ time. ∎
Statement 8.85.
The time complexitity of computing the values of the satellite data $counter_{E}$, and $counter_{F}$ for all rules in $R\langle E\cup F,\emptyset\rangle$ is $O(n)$.
Proof.
When Algorithm CAD computes the values of the satellite data $counter_{E}$ and $counter_{F}$ for all rules of the GTRS $R\langle E\cup F,\emptyset\rangle$, it carries out at most $2n$ additions, since $|C\langle E,F\rangle|\leq|W|\leq n$ and $|C\langle F,E\rangle|\leq|W|\leq n$. ∎
Statement 8.86.
The time complexity of constructing the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ is $O(n)$.
Proof.
We build the adjacency-list representation of the dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ by reading in search key order the nodes of the black-red tree $RBT6$ storing the relation $BSTEP\langle E\cup F,\emptyset\rangle$. Recall that for every ordered pair $({b},{c})\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component b{b}, the second component, c{c} is satellite data. Then the time complexity of building the adjacency-list representation of the dpwpa $AUX[E;F]$ is $O(|C\langle E\cup F,\emptyset\rangle|+|A[E;F]|)$. By Statement 4.28, $|C\langle E\cup F,\emptyset\rangle|\leq|W|\leq n$. By Statement 6.64, $|A[E;F]|=|BSTEP\langle E\cup F,\emptyset\rangle|\leq\mathrm{size}(E)+size(F)=2n$. Therefore the time complexity of building the adjacency-list representation of the dpwpa $AUX[E;F]$ is $O(n)$. ∎
Statement 8.78 implies the following fact.
Statement 8.87.
The time complexitity of computing the values of the attributes $equal_{E}$ and $equal_{F}$ of all constants ${b}\in C\langle E\cup F,\emptyset\rangle$ is $O(n)$.
Statement 8.88.
The time complexitity of computing the values of the attributes $keeps_{E}$ and $keeps_{F}$ of all constants ${b}\in C\langle E\cup F,\emptyset\rangle$ is $O(n\,\mathrm{log}\,n)$ time.
Proof.
We now estimate the number of operations when computing the values of the attributes ${b}.keeps_{E}$ for all ${b}\in C\langle E\cup F,\emptyset\rangle$. For each rewrite rule $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\in R\langle E\cup F,\emptyset\rangle$, Algorithm CAD carries out $m$ searches in the red-black tree $RBT3$ and less than or equal to $m$ multiplications. By Statement 4.46, $\mathrm{size}(R\langle E\cup F,\emptyset\rangle)\leq 2n$. In this way, when we compute the values of the attribute $keeps_{E}$ for all rewrite rules $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\in R\langle E\cup F,\emptyset\rangle$, the number of searches in the red-black tree $RBT3$ is less than equal to $n$. Consequently, by Proposition 2.1, the searches in the red-black tree $RBT3$ take $O(n\,\mathrm{log}\,n)$ time.
Furthermore, the number of multiplications is less than or equal to $\sum_{\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\in R\langle E\cup F,\emptyset\rangle}m\leq\mathrm{size}(R\langle E\cup F,\emptyset\rangle)\leq 2n$ by Statement 4.46. Thus, the number of multiplications is less than or equal to $2n$.
For each ${b}\in C\langle E\cup F,\emptyset\rangle$, the number of comparisons of nonnegative integers during computing the value of the attribute ${b}.keeps_{E}$ is equal to the number of rules in $R\langle E\cup F,\emptyset\rangle$ with right-hand side b{b}. Consequently, the number of comparisons of nonnegative integers during computing the values of the attributes ${b}.keeps_{E}$ for all ${b}\in C\langle E\cup F,\emptyset\rangle$ is equal to $|R\langle E\cup F,\emptyset\rangle|$. By Statement 4.46 $|R\langle E\cup F,\emptyset\rangle|\leq n$. Consequently the time complexity of computing for all ${b}\in C\langle E\cup F,\emptyset\rangle$, the values of the attribute ${b}.keeps_{E}$ is $O(n\,\mathrm{log}\,n)$. Similarly, the time complexity of computing for all ${b}\in C\langle E\cup F,\emptyset\rangle$, the values of the attribute ${b}.keeps_{F}$ is $O(n\,\mathrm{log}\,n)$. ∎
The Lemma follows from Statements 8.808.88. ∎